OpenStudy (anonymous):

Find the coordinates of the points of intersection of the circle with the y-axis. (x +5 )^2 + (y-4)^2 =169

OpenStudy (mertsj):

Let x = 0 if you want to find the y intercepts.

OpenStudy (anonymous):

how do you know what x equals?

OpenStudy (mertsj):

Replace x with 0 because if the point is on the y axis, its x coordinate is 0.

OpenStudy (mertsj):

\[(0+5)^2+(y-4)^2=169\]

OpenStudy (anonymous):

if it was asking for x axis, would it be 0 too?

OpenStudy (anonymous):

I got (0,16)

OpenStudy (mertsj):

\[25+(y-4)^2=169\] \[(y-4)^2=144\] \[y-4=12 or y-4=-12\] \[y=16 or y = -8\]

OpenStudy (mertsj):

I don't know what you mean "asking for the x axis" What does that mean?

OpenStudy (anonymous):

I mean like.. if a question like this one is asking for the x axis.. would it equal 0 like the y axis does?

OpenStudy (mertsj):

If you wanted to find the x intercepts you would replace y with 0.

OpenStudy (anonymous):

ok so my answer would be (0,16) correct?

OpenStudy (anonymous):

wait it says coordinates..

OpenStudy (mertsj):

Because every point on the x axis has a y coordinate of 0

OpenStudy (mertsj):

Did you not see my work?

OpenStudy (anonymous):

ohh no I'll check it out

OpenStudy (mertsj):

This circle crosses the y axis in TWO places.

OpenStudy (anonymous):

oh

OpenStudy (mertsj):

So (0,16) and (0,-8)

OpenStudy (anonymous):

But wouldnt we just change 16 to negative?

OpenStudy (anonymous):

how did you end up with -12

OpenStudy (anonymous):

in your steps

OpenStudy (mertsj):